please help me solve this problem:
a farmer and his son leave a barn at the same time and walk in opposite driections checking a fence line. the son walks at the speed of 3.5km/h and the farmer at 4.0km/h. how much time will have elapsed when the farmer and his son are 2.5km apart

Answer by nerdybill(5411)

a farmer and his son leave a barn at the same time and walk in opposite driections checking a fence line. the son walks at the speed of 3.5km/h and the farmer at 4.0km/h. how much time will have elapsed when the farmer and his son are 2.5km apart
.
Apply the distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Let t = elapsed time
then
3.5t + 4t = 2.5
7.5t = 2.5
t = 2.5/7.5
t = 25/75
t = 1/3 of an hour (or 20 minutes)